Home Obs Webcams
Weather Temp's Motors
Home Status OR Ins.
Weather Temp's Sat Im
Motors Obs Webcams
Find -tasks- here

SONG technical logbook page

Choose site
Logbook for Tenerife selected
Logbook entries for all projects are shown
Add to tasks?
Should someone look more into this or is it fixed?
Date of event
Who are you?
Hardware group
Action performed
[Maximum size = 20MB]
Send out notification e-mail? Messages will always be send to slack -> SONG-operations.

Search the log:
DateDescriptionPerson(s)Hardware GroupAction performedAttachmentEditSiteTask(?)
2023-05-31 Observations stopped due to serious telescope error Frank Grundahl Telescope Checked monitor.log and saw an error message about the az. Drives. Started Astelos, to see the error there, but decided it was above my level so did not perform any telescope actions. Closed the sideports and went back to bed, since the mirror covers and dome were already closed. - comments by Mads: After a complete system reboot of the telescope control system the telescope was again operational. None Edit Tenerife
2023-05-19 No light on detector from 02 UTC Mads Guider The slit guider connection was the issue. Reset the connection from kvmtenerife as root None Edit Tenerife
2023-05-14 Dirty mirrors causing less flux Mads All Washed M1, cleaned M9 and the lens before M9 None Edit Tenerife
2023-05-14 Slit motor slide was getting stuck since some time Mads Spectrograph Greased the spindle rod. Afterwards moved the motor 300 times without it got stuck None Edit Tenerife
2023-04-20 Observations did not start after sunset Frank Servers 1. changed status to abort for two time-critical obs. inserted for lam Ser. - did not help ---- Could not see any error messages in the log files. Tried to stop/start or_scheduler, and then monitor ... did not work. Then stop monitor and close mirror-cover, flap and dome with Astelos. Power is still on. It is probably something simple.... but I cannot figure it out :-) ... Comment from Mads: The conductor had crashed in Aarhus None Edit Tenerife
2023-04-07 Dust values were not updated Mads Weather station Restarted the dust collecting software and the weather station software. No values are updated again. Important since calima is seen on the web cameras. None Edit Tenerife
2023-03-21 Observations did not start after sunset Mads None Slony was not replicating observing requests to Tenerife because of the ongoing slony work... stopped slony on central, removed triggers for song-mtkent, song-central, song-tenerife and re-added for song-tenerife and restarted slony which made it replicate to tenerife again. None Edit Tenerife
2023-02-25 Weather station had issues after bad weather. Mads Weather station Power cycled the weather station control board which brought the WXT520 unit back but there is no response from the Boltwood cloud sensor. None Edit Tenerife 164
2023-02-12 Wire for lower hatch of dome was starting to get cut Antonio, Mads, Pere Dome Antonio replaced the wire. Show Edit Tenerife
2023-01-26 Slit motor was moved to position 6 and in software fixed there. Mads, Pere and Antonio Spectrograph Antonio and Pere opened the spectrograph and inspected the motor. Mads moved it remotely and when this was possible it was fixed at location 6 until a major operation to inspect, find cause and fix is planned and possilbe. None Edit Tenerife
2023-01-24 The slit motor is stuck again - This is a seriously bad issue!!! Mads Spectrograph Stopped the observations. None Edit Tenerife 161
2023-01-16 The slit motor is stuck - This is a seriously bad issue!!! Mads Spectrograph Tried checking the code to see if the issue was there. Then tried slowly to move the motor after reset of error state. Did not work. Then power cycle. Did not work -> out of operation None Edit Tenerife 160
2023-01-09 Dome was vignetting the pupil a bit Mads Dome Stopped the observations. Moved dome to 10 degrees N (past the home switch) and started the observations again. None Edit Tenerife
2023-01-09 Weather station was reporting raining from clear sky Mads Weather station Seemed like some rain indicator values was triggered and stayed triggered afterwards. Restarted the weather station software which solved the issue (for now). None Edit Tenerife
2023-01-08 Saving spectre to disk is slow.... takes about 13-15 seconds in the SaveAsFITS function of the Andor SDK Mads Servers This started occurring after the power failure on the 12th of december where the save to disk time increased by 2 seconds. On the 20th of december the save to disk time increased to 13-15 seconds. Only an issue in the SaveAsFITS function... strange.\n\rFor now a temporary solution is to read out the raw data and create data structure manually. This reduced time save to disk time to 2 seconds. None Edit Tenerife 159
2023-01-05 Calibrations did not finish in time when observations could start Mads Servers Stopped the observations and waited until calibrations had finished and then started the observations. Investigations (remote) needs to be performed to figure out the issue. None Edit Tenerife 157
2023-01-05 Late in the night the spectrograph slit motor reported not stopping. Mads Motors Stopped observations and stopped monitor and closed down. None Edit Tenerife 158
2023-01-03 Boltwood cloud sensor not responding... Mads Weather station Power cycled the weather station control board. No effect. Maybe cable has come loose, unit not working or power supply not working. Need on site assistance. None Edit Tenerife 155
2022-12-21 We will close down for x-mas from tonight until the 2nd of January Mads None Merry x-mas and happy new year! None Edit Tenerife
2022-12-18 Slit guider hanging from around 2:45 UTC Mads Guiders Discovered it in the morning. Reset the USB connection (remotely) and restarted the guide software. This made it work again. None Edit Tenerife
2022-12-17 OR script failing, message on missing calib files -> slit motor was reporting - still moving - Mads Motors Restarted the pst_server on hw and homed motor 6 first. The one that was failing. This worked. Then homed all and observations could continue. None Edit Tenerife
2022-12-12 Power failure at the observatory. A number of issues: main server did not boot up. Later no connection to the weather station board Mads All Managed to find the login credentials for ss5drac interface and boot up ss5n1. Then no connection to the weather station moxas indicating a failure in the multi socket power board. None Edit Tenerife 154
2022-11-29 Old ss3n1 server issues - preventing slit guiding. Server not booking up after reboot Mads, Antonio, Pere Slit guider Antonio helped to check connections and finally moved the USB plug to ss4n1 (kvmtenerife) on which the software was made run. None Edit Tenerife
2022-11-29 Guider hanging from 21 UTC - no observations since Mads Guider None Edit Tenerife
2022-11-28 Dust on windows and mirrors Mads, Antonio, Pere, Frank Telescope and coude Antonio cleaned window below M6. None Edit Tenerife